The following events took place for Digital Vibe Manufacturing Company during March, the first month of its operations as a producer of digital video monitors:

a. Purchased $137,800 of materials.
b. Used $94,100 of direct materials in production.
c. Incurred $184,800 of direct labor wages.
d. Incurred $212,720 of factory overhead.
e. Transferred $423,160 of work in process to finished goods.
f. Sold goods for $645,000.
g. Sold goods with a cost of $365,000.
h. Incurred $86,260 of selling expense.
i. Incurred $71,950 of administrative expense.

a. Using the information given, prepare the March income statement for Digital Vibe Manufacturing Company. Refer to the Amount Descriptions list provided for the exact wording of the answer choices for text entries.

Digital Vibe Manufacturing Company

Income Statement

For the Month Ended March 31

1

2

3

4

Operating expenses:

5

6

7

Total operating expenses

8

b. Using the information given, determine the inventory balances at the end of the first month of operations.

Materials
Work in process
Finished goods
Answer #1

1) Income statement

Sales 645000
Cost of goods sold 365000
Gross profit 280000
Selling and administrative expense
Selling expense 86260
Administrative expense 71950
Total Selling and administrative expense 158210
Net operating income 121790

Ending inventory balance

Material 137800-94100 = 43700
Work in process 94100+184800+212720-423160 = 68460
Finished goods 423160-365000 = $58160